# VoCo-LLaMA: Towards Vision Compression with Large Language Models

### v1

@@ Abstract
Vision-Language Models (VLMs) have achieved remarkable success in various multi-modal tasks, but they are often bottlenecked by the limited context window and high computational cost of processing high-resolution image inputs and videos. Vision compression can alleviate this problem by reducing the vision token count. Previous approaches compress vision tokens with external modules and force LLMs to understand the compressed ones, leading to visual information loss. However, the LLMs’ understanding paradigm of vision tokens is not fully utilised in the compression learning process. We propose VoCo-LLaMA, the first approach to compress vision tokens using LLMs. By introducing Vision Compression tokens during the vision instruction tuning phase and leveraging attention distillation, our method distill how LLMs comprehend vision tokens into their processing of VoCo tokens. VoCo-LLaMA facilitates effective vision compression and improves the computational efficiency during the inference stage. Specifically, our method achieves minimal performance loss with a compression ratio of 576, resulting in up to 94.8% fewer FLOPs and 69.6% acceleration in inference time. Furthermore, through continuous training using time-series compressed token sequences of video frames, VoCo-LLaMA demonstrates the ability to understand temporal correlations, outperforming previous methods on popular video question-answering benchmarks. Our approach presents a promising way to unlock the full potential of VLMs’ contextual window, enabling more scalable multi-modal applications. The project page, along with the associated code, can be accessed via the following link.
@@ 1 Introduction
The advent of visual-language models [1, 2, 3, 4, 5, 6, 7, 8, 9, 10] has led to significant advancements in visual understanding. Particularly, high-resolution image encoding [7, 8] and the incorporation of more video frames [9, 10] have been shown to enhance the capabilities of both large visual-language models and large video-language models, respectively. However, the large number of vision tokens occupies a substantial portion of the valuable context window of the large language model, leading to expensive computational costs, as shown in Figure 1 (a). For instance, when using high-resolution image inputs in LLaVA-1.6 [7], a single image with a resolution of 672 672 is divided into four smaller patches, each encoded with a 336 336 resolution input. This process yields an image representation consisting of 2304 vision tokens, occupying over half of the context length. Furthermore, as the number of input images increases, the context window for text will be further bottlenecked. For example, Vicuna-1.5 [11] can only handle approximately 7 frames (7 576 = 4032 tokens) within its 4k context length, given away the text inputs. [9, 10] investigate the efficacy of extending the context length to the million-level mark to mitigate this issue, but this approach necessitates expensive computational resources (e.g., [9] requires over 1000 v4 TPUs) and engineering efforts in data preparation and framework development.
Effective representation of visual information within limited context length is crucial for addressing the challenges. A promising approach is to reduce the number of vision tokens representing an image, ideally to just a handful or even a single token. As illustrated in Figure 1 (b), previous methods [1, 3, 4, 6, 12, 13] have exploited the use of Q-Former [1] or straight-forward pooling operations to "compress" the encoded vision tokens. These methods compress the vision tokens with external modules and force LLMs to understand the compressed ones. This may disrupt the established paradigm of how LLMs comprehend original uncompressed vision tokens and bring loss of visual information. However, LLMs with uncompressed vision token input have already possessed visual understanding capabilities. Therefore, a more intuitive approach of vision compression is to distill how LLMs understand vision tokens into their processing of compressed tokens.
@@
We propose VoCo-LLaMA, the first vision compression method that exploits the inherent functionality of large language models. As demonstrated in Figure 1 (c), we introduce Vision Compression (VoCo) tokens between visual and text tokens. By simply modifying the attention mechanism, we first isolate the visual and text tokens by inserting VoCo tokens. Subsequently, we establish an exclusive interaction pathway between the visual and text tokens via VoCo tokens. This facilitates the LLM itself to compress and distill the parsing vision tokens, specifically the transformer activations on top of them, into compact VoCo tokens. Benefiting from this compression strategy, our method achieves modality alignment without the specialized text-vision cross-modal fusion module design [1, 12].
Our method can learn vision compression through standard visual instruction tuning, eliminating the need for incurring additional training phase. Furthermore, when handling multiple repetitive visual inputs in the inference stage, VoCo-LLaMA enables caching and reusing compressed transformer activations, improving computational efficiency and reducing storage requirements for caching. VoCo-LLaMA can achieve a 576 compression rate (compressing 576 vision tokens into a single VoCo token) while maintaining 83.7% performance. Additionally, we observe a reduction in inference computation costs, with up to 99.8% in cache storage, 94.8% FLOPs and 69.6% in inference time.
To further investigate the efficacy of our proposed method in handling video input, we utilize the time-series VoCo token sequence obtained by compressing video frames to further explore the temporal modeling capabilities of VoCo-LLaMA. Without any additional designs, VoCo-LLaMA outperforms existing vision compression methods on common video question-answering benchmarks, with each video frame compressed into an equal number of tokens.
@@ 2 Related Work
LLMs and Text Compression. In recent years, large language models (LLMs) have sparked a technological revolution. As the scale of training data and model size continue to expand, models [11, 14, 15, 16, 17, 18, 19, 20] have demonstrated exceptional capabilities in understanding and generating language, showcasing the powerful abilities of LLMs in handling complex language tasks. In particular, models such as the LLaMA series [17, 18, 20], Vicuna [11], and Mistral [19] have emerged as foundational models or main components in many research works. However, the limited context window size in LLMs has long been a widely discussed topic in the natural language processing field. Compressing text to conserve valuable window length has been proven to be an efficient approach. Long-standing works, including [21, 22, 23, 24, 25], focus on storing text representations in transformers to achieve dense information representation. [26, 27] have demonstrated the effectiveness of distilling long text information into prompt-free student models. [28] has made significant progress in compressing information in long contexts by learning soft prompts. In a similar vein, recent studies [29, 30, 31] have explored the potential applications of compressing text in large language models. Although the context length limitation is a likewise pervasive challenge in vision-language models, the discussion of compressing visual information has been relatively understudied compared to the language model domain. Our work pioneers the use of LLMs’ learning capabilities to compress and distill information in vision tokens, aiming to bridging this gap in the field of VLMs.
@@
VLMs and Vision Compression. The success of LLMs has inspired significant progress in vision language models (VLMs). By integrating visual encoders with LLMs, VLMs can effectively achieve cross-modal understanding and other visual language tasks through modality alignment and instruction tuning. Previous methods [1, 2, 3, 4, 5, 6, 7, 8, 13, 32] have substantiated the success of this training paradigm in visual understanding. The successful application of VLMs on images has also been rapidly extended to the video domain [9, 10, 12, 33, 34, 35, 36, 37, 38, 39]. With the input of higher-resolution images [7, 8] and more video frames [9, 10], VLMs can capture rich visual information for understanding and generation. However, as the number of vision tokens representing an input image increases, they take up a significant portion of the limited context window of language models, and can even exceed it. To address this, previous methods [1, 3, 4, 6, 13] have largely employed Q-Former [1], which maps images to fixed-length tokens in the language embedding space through learnable queries, compressing visual information. A more recent approach [12] has applied average pooling with a learnable linear layer to compress visual features through multi-stage training strategy. Although these methods perform moderately well at lower compression multiples, they cause a significant loss of valuable visual information when the number of compressed tokens reduces. VoCo-LLaMA distills the approach of LLMs in understanding vision tokens into their processing of compressed tokens, significantly reducing information loss during the vision compression process.
@@ 3.1 Vision Compression
Given a paired image and text input, we follow the design of most vision-language models (VLMs) and encode the image into a sequence of vision tokens , where is the number of the output patches from the visual encoder. Similarly, the text input is encoded into a sequence of text tokens . Consider an original, unmodified standard vision-language model (denoted as ), exemplified by LLaVA [2], depicted in Figure 2 (a). During visual instruction tuning, leverages both vision tokens and text tokens to predict the output through self-regression, and learns the distribution . For image compression models, our goal is to employ a compact set of compressed tokens to efficiently represent the sparse vision token set . Additionally, we aim to generate outputs that closely approximates the outputs of the original model when presented with identical image and text inputs.
Specifically speaking, with an image encoded as vision tokens , we formulate the image compression distillation process as learning a compression model that generates the output conditioned on the compressed tokens and the text tokens . This is achieved by learning a probability distribution . The optimization objective of is to minimize the loss function:
@@ 3.2 VoCo-LLaMA
As illustrated in Figure 2 (b), we propose VoCo-LLaMA, which introduces special Vision Compression (VoCo) tokens to leverage the LLM’s ability to compress and understand the compression representations of images. The input sequence to the large language model is formed by concatenating the vision tokens, the special tokens, and the text tokens, yielding a sequence:
In the training phase, we employ a two-stage attention mechanism. Initially, we impose a constraint on the text tokens, explicitly preventing them from attending to the original vision tokens, while forcing them to exclusively attend to the compressed and distilled VoCo tokens. Subsequently, the vision tokens are subjected to continuous attention from the VoCo tokens due to the casual transformer. This deliberate design ensures that the text tokens solely capture the distilled visual information encoded in the VoCo tokens, rather than directly interacting with the original vision tokens, achieving compressing distillation from vision tokens to compressed tokens.
The compression process of VoCo-LLaMA can be elegantly implemented by strategically modifying the attention mask. Specifically, we set the attention weights between the text tokens and the vision tokens to , effectively rendering the text tokens "isolated" to the vision tokens. Formally, let denote the attention mask, where if token attends to token , and otherwise. We define the modified attention mask as:
It is worth noting that our modification to the attention mask is based on the casual attention mask implementation in decoder-only models (e.g. LLaMA [17]), which is initialized as a lower-left triangular matrix. In practice, VoCo-LLaMA can be effectively trained under the standard supervised fine-tuning paradigm, leveraging the abundant image-text data readily available in VLMs. Furthermore, the VoCo token can be compactly represented as a set of transformer activations, allowing them to be cached to enhance inference efficiency.
VoCo-LLaMA enables the large language models to learn the compression process of vision tokens, , while simultaneously learning to understand the compressed VoCo tokens, i.e., learning the distribution . With VoCo-LLaMA, the optimization objective in Equation 2 can be further defined as:
Moreover, our design affords an additional advantage: the VoCo tokens exclusively interact with the input text during the compression process, establishing a seamless pathway between the visual and text modalities. Consequently, this enables us to effortlessly align the compressed visual information with the text modality, eliminating the need for the specialized text-vision aware cross-modal alignment module design[1, 12].
@@ 3.3 Temporal Modeling
When giving a sequence of video frames and a corresponding text input, VoCo-LLaMA, as introduced in Section 3.2, could serve as a vision encoding compressor to encode each frame. As shown in Figure 3, we obtain the compressed representation for each frame with the video frame tokens and the compression token . This yields a sequence of compressed video tokens, denoted by .
By harnessing the current capabilities of VoCo-LLaMA to interpret tokens derived from compressing individual frames, we can further enhance its functionality to handle temporal inputs. Specifically, we adopt a continue training process on VoCo-LLaMA by utilizing the time-series compressed token sequences , enabling it to capture and comprehend the temporal correlations among the compressed video tokens. With the inclusion of text tokens , VoCo-LLaMA learns the distribution . This training strategy builds upon existing image compression capabilities and allows VoCo-LLaMA to focus on temporal modeling, thereby streamlining the video understanding process.
@@ 3.4 Implementation Details
Regarding the training strategy and data, as mentioned earlier in Section 3.2, VoCo-LLaMA only requires learning to insert and compress VoCo tokens during the vision instruction tuning stage. We follow the common VLMs [2, 7] to encode the image input into vision tokens with vision encoder and a linear projector. We employ the pre-trained CLIP-ViT-L [40] as our visual encoder. For pre-trained large language models, we utilize Vicuna-7B [11]. Without introducing VoCo tokens, we first align the visual encoder and language model using the LLaVA-filtered CC3M [41] dataset with visual encoder and language model keeping frozen. During the instruction tuning phase of VoCo-LLaMA, incorporating multiple image understanding tasks is crucial for learning a scalable image compression model. Therefore, we construct the instruction pairs inspired by [12] using [7]. For video tuning, we further utilize WebVid [42] and the QA-pairs of Video-ChatGPT [39].
We conducted experiments on several common compression strategies with the same training setting and data for comparison. For the compression strategy with Q-Former, we employ the architecture in [1] and configure the query number to one, resulting in a single compression token. For the compression strategy with average pooling, we follow the design of the single content token in [12], as illustrated in Figure 1 (b). For more details on the implementation, please refer to the Appendix A.
@@ 4 Experiments
4.1 Datasets
In this work, we conduct experiments on several common visual understanding benchmarks for vision compression. In particular, we report results on GQA [43], MMB (MMBench) [44], MME [45], POPE [46], SEED-Bench [47], SQAI (Image-based setting in ScienceQA) [48] and VQAv2 (VQA V2) [49]. By observing the model’s performance on these image understanding benchmarks before and after compression (i.e. with initial vision tokens / VoCo tokens), we can observe the effects of the visual information loss that occurs during the vision compression process. We evaluate the performance on these visual understanding benchmarks in accordance with the details outlined in [2]. As for the video domain, we evaluate the zero-shot performance on several video question-answering benchmarks. MSVD-QA [50] is a video QA dataset consisting of 1,970 video clips with 50,505 QA pairs, built upon the Microsoft Research Video Description Corpus [51]. MSRVTT-QA [50] is a large-scale video QA dataset featuring 10K videos and 243K QA pairs with complex scenes, based on the MSR-VTT dataset [52]. ActivityNet-QA [53] is a fully annotated video QA dataset containing 58K QA pairs derived from 5,800 complex web videos from the ActivityNet dataset [54].
4.2 Vision Compression Configuration
4.3 Results
Vision Compression. Table 1 presents the results of VoCo-LLaMA on vision compression. To explore the maximum potential of our method, we report the highest achievable compression ratio, which compresses vision tokens into a single VoCo token. We report results of our compression model on various common visual understanding benchmarks, as well as the compression retention rates defined based on upper and lower bound models introduced in Section 4.2. It can be observed that our method preserves the original visual information to a large extent, even at an extremely high compression ratio of 576. Specifically, we achieved an average compression retention rate of 83.7% across seven widely used benchmarks. Especially on MMBench, POPE and VQAv2, our method retained more than 85% of the performance during compression. The results indicate that VoCo-LLaMA can effectively compress vision tokens. Moreover, our method consistently outperforms the performance lower bound model of random compression across all benchmarks. This demonstrates that the advantages of VoCo-LLaMA, such as significant reductions in context length and improved calculation efficiency, outweigh any potential performance loss caused by compression.
We additionally compare our method with previous common learning-based approaches (i.e., Q-Former and average pooling) for vision token compression. Our method significantly outperforms previous methods across all benchmarks. Specifically, we observe an improvement of 19.6% in average compression retention rate, surpassing the average pooling compression strategy. In contrast, while Q-Former has demonstrated impressive capabilities in capturing visual features with 32 queries, its performance undergoes a substantial decline when the query count is reduced to a single digit. This proves that the compression method of VoCo-LLaMA, which utilizes the knowledge distillation from large language models itself, maintains more valuable vision information than that of average pooling or query-based compression.
@@ 4.3 Results
Number of VoCo tokens. We evaluate the impact of the number of VoCo tokens on vision compression performance. As shown in Table 2, we report the results of VoCo-LLaMA on four widely used benchmarks as the number of VoCo tokens increases from 1 to 128, along with the average compression retention rates across four benchmarks. Figure 4 illustrates the trend of compression performance retention as the number of VoCo tokens varies, where the green and red lines represent the upper and lower bounds of compression performance, respectively. We adopted the same training settings and data as in the main experiments.

### v2

@@ Abstract
Vision-Language Models (VLMs) have achieved remarkable success in various multi-modal tasks, but they are often bottlenecked by the limited context window and high computational cost of processing high-resolution image inputs and videos. Vision compression can alleviate this problem by reducing the vision token count. Previous approaches compress vision tokens with external modules and force LLMs to understand the compressed ones, leading to visual information loss. However, the LLMs’ understanding paradigm of vision tokens is not fully utilised in the compression learning process. We propose VoCo-LLaMA, the first approach to compress vision tokens using LLMs. By introducing Vision Compression tokens during the vision instruction tuning phase and leveraging attention distillation, our method distill how LLMs comprehend vision tokens into their processing of VoCo tokens. VoCo-LLaMA facilitates effective vision compression and improves the computational efficiency during the inference stage. Specifically, our method can achieve a 576 compression rate while maintaining 83.7% performance. Furthermore, through continuous training using time-series compressed token sequences of video frames, VoCo-LLaMA demonstrates the ability to understand temporal correlations, outperforming previous methods on popular video question-answering benchmarks. Our approach presents a promising way to unlock the full potential of VLMs’ contextual window, enabling more scalable multi-modal applications.
@@ 1 Introduction
The advent of visual-language models [25, 30, 60, 13, 3, 59, 29, 5, 32, 47] has led to significant advancements in visual understanding. Particularly, high-resolution image encoding [29, 5] and the incorporation of more video frames [32, 47] have been shown to enhance the capabilities of both large visual-language models and large video-language models, respectively. However, the large number of vision tokens occupies a substantial portion of the valuable context window of the large language model, leading to expensive computational costs. For instance, when using high-resolution image inputs in LLaVA-1.6 [29], a single image with a resolution of 672 672 is divided into smaller patches, each encoded with a 336 336 resolution input. This process yields an image representation consisting of 2880 vision tokens, occupying over half of the context length. As the number of input images increases, the context window for text will be further bottlenecked. [32, 47] investigate the efficacy of extending the context length to the million-level mark to mitigate this issue, but this approach necessitates expensive computational resources (e.g., [32] requires over 1000 v4 TPUs) and engineering efforts in data and framework development.
To address this issue, previous methods [25, 60, 13, 59, 28, 11] have exploited Q-Former [25] or Re-sampler [1] to “compress” the encoded vision tokens. As illustrated in Fig. 1 (a), these kind of methods compress the vision tokens with external modules and force LLMs to understand the compressed ones.
@@
Given that the LLM can effectively understand uncompressed vision tokens [31], it has great potential to perform token compression on its own. Therefore, we propose VoCo-LLaMA, the first vision compression method that leverages the inherent capabilities of large language models to our best knowledge. As demonstrated in Fig. 1 (b), we introduce Vision Compression (VoCo) tokens between visual and text tokens. By modifying the attention mechanism, we ensure that VoCo tokens attend exclusively to visual tokens, while text tokens attend solely to VoCo tokens. Subsequently, we establish an exclusive interaction pathway between the visual and text tokens via VoCo tokens. This facilitates the LLM itself to compress and distill the parsing vision tokens, specifically the transformer activations on top of them, into compact VoCo tokens.
Building upon this, we further investigate the efficacy of VoCo-LLaMA in handling video input. The total number of visual tokens for each video can be substantial, far exceeding the context length of large language models (LLMs), making it impractical to utilize VoCo-LLaMA to compress all the tokens simultaneously. To address this issue, we first employ VoCo-LLaMA to compress the visual tokens of each frame into voco tokens. These voco tokens are subsequently concatenated to form a sequential token series. VoCo-LLaMA then extracts both visual and temporal information from this series to facilitate video understanding tasks. With this effective design, the LLMs can handle much longer videos within the same context length.
During inference, VoCo-LLaMA mitigates the issue of limited context length in LLM by employing a two-stage forward process. The fist stage compresses visual tokens of each image into a reduced set of VoCo tokens, while the second stage completes the task by utilizing both VoCo tokens and text tokens. Moreover, VoCo tokens can be cached and reused when handling various tasks involving identical visual inputs, thereby enhancing computational efficiency and reducing storage requirements compared to maintaining the entire KV-cache for uncompressed visual tokens. Experimental results on various benchmarks demonstrate that VoCo-LLaMA achieves a 576x compression rate while maintaining 83.7% of the original performance. Additionally, significant reductions in inference computation costs were observed, including up to 99.8% in cache storage, 94.8% in FLOPs, and 69.6% in inference time.
Our core contributions are summarized as follows:
We propose VoCo-LLaMA, the first approach to compress vision tokens by leveraging the inherent capabilities of large language models, thereby eliminating the need for any external modules.
We extend VoCo-LLaMA from image input to video input, which allows the LLM to handle approximately 200 times more video frames while maintaining its video understanding capabilities.
Extensive experiments on image and video benchmarks demonstrate the effectiveness of our method, showcasing superior performance in both token compression and inference efficiency compare to various existing baselines.
@@ 2 Related Work
LLMs and Text Compression. In recent years, large language models (LLMs) have sparked a technological revolution. As the scale of training data and model size continue to expand, models [10, 6, 41, 51, 48, 49, 22, 20] have demonstrated exceptional capabilities in understanding and generating language. In particular, models such as the LLaMA series [48, 49, 20, 10] have emerged as foundational models or main components in many research works. However, the limited context window size in LLMs has long been a widely discussed topic. Text compression has been proven to be an efficient approach. Long-standing works, including [12, 33, 44, 52, 57], focus on storing text representations in transformers to achieve dense information representation. [2, 46] have demonstrated the effectiveness of distilling long text information into prompt-free student models. In a similar vein, recent studies [50, 9, 40, 16] have explored the potential applications of compressing text in large language models. However, the discussion of compressing visual information has been relatively understudied compared to the language model domain. Our work pioneers the use of LLMs’ learning capabilities to compress vision information, aiming to bridging this gap in the field of VLMs.
@@
VLMs and Vision Compression. The success of LLMs has inspired significant progress in vision language models (VLMs). By integrating visual encoders with LLMs, VLMs can effectively achieve cross-modal understanding through instruction tuning. Previous methods [25, 30, 60, 13, 3, 59, 29, 5, 11, 1] have substantiated the success of this training paradigm in visual understanding. The successful application of VLMs on images has also been rapidly extended to the video domain [32, 47, 28, 19, 23, 58, 26, 34, 37, 39]. With the input of higher-resolution images [29, 5] and more video frames [32, 47], VLMs can capture rich visual information. However, as the number of vision tokens representing an input image increases, they take up a significant portion of the limited context window of language models, and can even exceed it. To address this, previous methods [25, 60, 13, 59, 11] have largely employed Q-Former [25], which maps images to fixed-length tokens in the language embedding space through learnable queries, compressing visual information. A more recent approach [28] has applied average pooling with a learnable linear layer to compress visual features through multi-stage training strategy. Although these methods perform moderately well at lower compression multiples, they cause a significant loss of valuable visual information when the number of compressed tokens reduces. VoCo-LLaMA distills the approach of LLMs in understanding vision tokens into their processing of compressed tokens, significantly reducing information loss during the vision compression process.
@@ 3.1 Vision Compression
Given a paired image and text input, we follow the design of most vision-language models (VLMs) and encode the image into a sequence of vision tokens , where is the number of the output patches from the visual encoder. Similarly, the text input is encoded into a sequence of text tokens . Consider an original, unmodified standard large vision language model (denoted as ), exemplified by LLaVA [30], depicted in Fig. 2 (a). During visual instruction tuning, leverages both vision tokens and text tokens to predict the output , and learns the distribution . For image compression models, our goal is to employ a compact set of compressed tokens to efficiently represent the vision token set . Additionally, we aim to generate outputs that closely approximates the outputs of the original model when presented with identical image and text inputs.
With an image encoded as vision tokens , we formulate the image compression distillation process as learning a compression model that generates the output conditioned on the compressed tokens and the text tokens . This is achieved by learning a probability distribution . The optimization objective of is to minimize the loss function:
@@ 3.2 VoCo-LLaMA
As illustrated in Fig. 2 (b), VoCo-LLaMA leverages the LLM’s ability to compress visual tokens into compact Vision Compression (VoCo) tokens and learns to understand image through these VoCo tokens. The input sequence to the large language model is formed by concatenating the vision tokens, the special tokens, and the text tokens, yielding a sequence:
In the training phase, we employ a two-stage attention mechanism. Initially, we impose a constraint on the text tokens, explicitly preventing them from attending to the original vision tokens, and requiring them to exclusively attend to the compressed and distilled VoCo tokens. Subsequently, the vision tokens are subjected to continuous attention from the VoCo tokens due to the casual attention mechanism. This deliberate design ensures that the text tokens solely capture the distilled visual information encoded in the VoCo tokens, rather than directly interacting with the original vision tokens, thereby achieving effective compression from vision tokens to compressed tokens.
The compression process of VoCo-LLaMA can be elegantly implemented by modifying the attention mask. Specifically, we set the attention weights between the text tokens and the vision tokens to , effectively rendering the text tokens “isolated” to the vision tokens. Formally, let denote the attention mask, where if token attends to token , and otherwise. We define the attention mask as:
In practice, VoCo-LLaMA can be effectively trained under the standard supervised fine-tuning paradigm, leveraging the abundant image-text data readily available in VLMs. Furthermore, the VoCo token can be compactly represented as a set of Transformer activations, allowing them to be cached to enhance inference efficiency, which will be discussed in Sec. 3.3.
VoCo-LLaMA enables the large language models to learn the compression process of vision tokens, , while simultaneously learning to understand the compressed VoCo tokens. We define the target learning distribution as follows:
the optimization objective in Eq. 1 can be defined as:
3.3 Reuse of VoCo Cache
During inference, VoCo-LLaMA mitigates the issue of limited context window size by dividing the single forward pass into two phases. As illustrated in Fig. 4, the first forward pass takes [vision tokens, VoCo tokens] as input to compress visual information into Transformer activations upon VoCo tokens. The second forward pass takes [text tokens] as input and load VoCo activations as KV Cache.
Moreover, VoCo tokens derived from the first forward pass can be cached and reused when handling various tasks involving identical image/video inputs, thereby enhancing computational efficiency and reducing storage requirements compared to maintaining the entire KV-cache for uncompressed visual tokens. For more details on the inference implementation, please refer to the supplementary material.
@@ 3.4 Temporal Modeling
When giving a sequence of video frames and a corresponding text input, the token length for the entire video far exceeds LLM context length, as shown in Fig. 3 (a). To solve this issue, VoCo-LLaMA divides video input into smaller segments and input these segments into LLM with VoCo tokens . As shown in Fig. 3 (b), all the frames are compressed into VoCo activations. Specifically, we obtain the compressed representation for video segment tokens through . This yields a sequence of KV Cache representing compressed video tokens, denoted by .
Having obtained the time-series compressed cache sequences , we enable language model to capture and comprehend the temporal correlations among the compressed video tokens. With the inclusion of text tokens , VoCo-LLaMA learns the distribution . We adopt a continue training process based on VoCo-LLaMA with image compression capabilities, allows the model to focus on temporal modeling, thereby streamlining the video understanding process.
@@ 3.5 Implementation Details
Regarding the training strategy and data, as mentioned earlier in Sec. 3.2, VoCo-LLaMA only requires learning to insert and compress VoCo tokens during the vision instruction tuning stage. We follow the common VLMs [30, 29] to encode the image input into vision tokens with vision encoder and a linear projector. We employ the pre-trained CLIP-ViT-L [43] as our visual encoder. For pre-trained large language models, we utilize Vicuna-7B [10]. Without introducing VoCo tokens, we first align the visual encoder and language model using the LLaVA-filtered CC3M [45] dataset with visual encoder and language model keeping frozen. During the instruction tuning phase of VoCo-LLaMA, incorporating multiple image understanding tasks is crucial for learning a scalable image compression model. Therefore, we construct the instruction pairs inspired by [28] using [29]. For video tuning, we further utilize WebVid [4] and the QA-pairs of Video-ChatGPT [39]. Moreover, gradient checkpointing strategies are employed to reduce computational cost during training.
We conducted experiments on several common compression strategies with the same training setting and data for comparison. For the compression strategy with Q-Former, we employ the architecture in [25] and configure the query number to one, resulting in a single compression token. For the compression strategy with average pooling, we follow the design of the single content token in [28]. For more details on the training and inference implementation, please refer to the supplementary material.
@@ 4 Experiments
4.1 Datasets
In this work, we conduct experiments on several common visual understanding benchmarks for vision compression. In particular, we report results on GQA [21], MMB (MMBench) [35], MME [15], POPE [27], SEED-Bench [24], SQAI (Image-based setting in ScienceQA) [36] and VQAv2 (VQA V2) [17]. By observing the model’s performance on these image understanding benchmarks before and after compression (i.e. with initial vision tokens / VoCo tokens), we can observe the effects of the visual information loss that occurs during the vision compression process. We evaluate the performance on these visual understanding benchmarks in accordance with the details outlined in [30]. As for the video domain, we evaluate the zero-shot performance on several video question-answering benchmarks. MSVD-QA [53] is a video QA dataset consisting of 1,970 video clips with 50,505 QA pairs, built upon the Microsoft Research Video Description Corpus [7]. MSRVTT-QA [53] is a large-scale video QA dataset featuring 10K videos and 243K question-answering pairs with complex scenes, based on the MSR-VTT dataset [54]. ActivityNet-QA [56] is a fully annotated video QA dataset containing 58K question-answering pairs derived from 5,800 complex web videos from the ActivityNet dataset [18].
4.2 Vision Compression Configuration
4.3 Results
Vision Compression. Tab. 1 presents the results of VoCo-LLaMA on vision compression. To explore the maximum potential of our method, we report the highest achievable compression ratio, which compresses vision tokens into one single VoCo token. We report results of our compression model on various common visual understanding benchmarks, as well as the compression retention rates defined based on upper and lower bound models introduced in Sec. 4.2. It can be observed that our method preserves the original visual information to a large extent, even at an extremely high compression ratio of 576. Specifically, we achieved an average compression retention rate of 83.7% across seven widely used benchmarks. Especially on MMBench, POPE and VQAv2, our method retained more than 85% of the performance during compression. The results indicate that VoCo-LLaMA can effectively compress vision tokens. Moreover, our method consistently outperforms the performance lower bound model of random compression across all benchmarks. This demonstrates that the advantages of VoCo-LLaMA, such as significant reductions in context length and improved calculation efficiency, outweigh the performance loss caused by compression.
| Method | N | GQA | POPE | SQAI | VQAT |
| LLaMA-VID [28] | 16 | 58.2 | 83.1 | 67.4 | 50.8 |
| 4 | 56.2 | 83.5 | 68.7 | 49.1 | |
| 1 | 55.5 | 83.1 | 68.8 | 49.0 | |
| VoCo-LLaMA | 1 | 58.3 | 85.0 | 69.5 | 52.7 |
| Method | Token | MMB | GQA | VQAv2 | SEED |
| VoCo-LLaMA | 32 | 59.4 | 60.2 | 75.3 | 56.2 |
| 16 | 58.3 | 58.9 | 74.9 | 55.8 | |
| 4 | 59.7 | 58.0 | 73.5 | 55.2 | |
| 1 | 57.9 | 56.1 | 71.2 | 53.0 |
| Method | Token | RefCOCO | RefCOCO+ | RefCOCOg | GRIT | Avg. | |||||
| val | test A | test B | val | test A | test B | val (U) | test (U) | refexp | |||
| Upper Bound | 256 | 87.01 | 90.61 | 80.24 | 81.60 | 87.36 | 72.12 | 82.27 | 82.19 | 69.34 | 100% |
| VoCo-LLaMA | 8 | 85.17 | 88.92 | 79.21 | 80.02 | 85.13 | 70.22 | 80.36 | 80.64 | 68.59 | 90.7% |
| 1 | 83.29 | 86.89 | 77.87 | 77.62 | 83.02 | 67.74 | 78.32 | 78.06 | 67.69 | 79.9% | |
| Lower Bound | 1 | 68.34 | 72.96 | 68.03 | 62.58 | 64.77 | 50.65 | 62.30 | 62.99 | 60.50 | 0% |
| Method | Token | RefCOCO | RefCOCO+ | RefCOCOg | Avg. | |||||
| val | test A | test B | val | test A | test B | val (U) | test (U) | |||
| Upper Bound | 256 | 75.61 | 44.26 | 104.83 | 56.42 | 40.98 | 68.25 | 62.71 | 65.58 | 100% |
| VoCo-LLaMA | 8 | 73.87 | 43.13 | 102.71 | 55.34 | 39.91 | 67.00 | 61.59 | 64.45 | 91.3% |
| 1 | 71.92 | 41.81 | 94.50 | 53.98 | 38.96 | 65.35 | 60.46 | 63.17 | 81.6% | |
| Lower Bound | 1 | 56.73 | 31.82 | 78.09 | 43.71 | 30.26 | 52.22 | 50.49 | 53.22 | 0% |
We additionally compare our method with previous common learning-based approaches (i.e., Q-Former and average pooling) for vision token compression. Our method significantly outperforms previous methods across all benchmarks. Specifically, we observe an improvement of 19.6% in average compression retention rate, surpassing the average pooling compression strategy. In contrast, while Q-Former has demonstrated impressive capabilities in capturing visual features with 32 queries, its performance undergoes a substantial decline when the query count is reduced to a single digit. This proves that our VoCo-LLaMA, which utilizes the knowledge distillation from language models itself, maintains more valuable vision information than that of average pooling or query-based compression.